# Distilling Implicit Multimodal Knowledge into LLMs for Zero-Resource Dialogue Generation

### v1

@@ V-B Human Evaluation Results
TABLE III: Ablation study. Zero Data means in a fully zero-resource scenario.
@@
V-C Ablation Study
To evaluate the individual contributions of the components within VIKDF, we conduct an ablation study by sequentially removing each component and assessing the impact on performance metrics (BLEU-1 and ROUGE-L) across both Reddit Conversation and Image-Chat datasets. In the absence of BVIF, we adopt the vision-to-language generative learning as outlined in [7]. The results, detailed in Table III, demonstrate a consistent decline in performance upon the exclusion of any component, underscoring their collective importance to the framework’s efficacy. Notably, the exclusion of TIM yields the most pronounced decline in performance metrics. This indicates TIM’s critical function in maintaining alignment between textual and visual modalities, a foundational aspect of generating coherent and contextually relevant dialogues. Additionally, omitting BVIF leads to a noticeable dip in performance metrics. This underscores BVIF’s importance in seamlessly integrating distilled visual knowledge into the dialogue generation process, further enhancing the model’s ability to produce contextually rich and engaging dialogues. The removal of TAMIM and IAMTM also leads to decreased performance. This result highlights their significance in enriching the model’s capability to infer and align multimodal knowledge, thereby facilitating a more nuanced dialogue generation process.
((a))
((b))
@@ V-C Ablation Study
Figure 7: Example of visual implicit knowledge textualized by VIKDF.
V-D Case study
To provide a more detailed analysis of visual implicit knowledge, we attempt to textualize the visual implicit knowledge through LLMs, following the process illustrated in Figure 7. The process involves transforming the dialogue context through IQ-Former that distills visual implicit knowledge, subsequently textualizing it via an LLM by a text prompt. We can see that the textual output illustrates the model’s capacity to encapsulate complex human emotions and curiosities about the sky. This demonstrates the model’s adeptness at integrating and expressing visual implicit knowledge.
V-E Impact of Query Vector Quantity
@@
VI Conclusion and Future Work
@@
Acknowledgments
This research is supported by the National Natural Science Foundation of China (No. 62006034).

### v2

@@ 5.2 Human Evaluation Results
Table 3: Ablation study. Zero Data means in a fully zero-resource scenario.
@@
5.3 Ablation Study
To evaluate the individual contributions of the components within VIKDF, we conduct an ablation study by sequentially removing each component and assessing the impact on performance metrics (BLEU-1 and ROUGE-L) across both Reddit Conversation and Image-Chat datasets. In the absence of BVIF, we adopt the vision-to-language generative learning as outlined in [8]. The results, detailed in Table 3, demonstrate a consistent decline in performance upon the exclusion of any component, underscoring their collective importance to the framework’s efficacy. Notably, the exclusion of TIM yields the most pronounced decline in performance metrics. This indicates TIM’s critical function in maintaining alignment between textual and visual modalities, a foundational aspect of generating coherent and contextually relevant dialogues. Additionally, omitting BVIF leads to a noticeable dip in performance metrics. This underscores BVIF’s importance in seamlessly integrating distilled visual knowledge into the dialogue generation process, further enhancing the model’s ability to produce contextually rich and engaging dialogues. The removal of TAMIM and IAMTM also leads to decreased performance. This result highlights their significance in enriching the model’s capability to infer and align multimodal knowledge, thereby facilitating a more nuanced dialogue generation process.
Table 4: Performance comparison of VIKDF integrated with different LLMs.
Methods Reddit Conversation Image-Chat (Zero Data)
BLEU-1 ROUGE-L BLEU-1 ROUGE-L
LlamaLoRA 14.81 13.13 13.34 12.04
QwenLoRA 15.04 13.26 13.44 12.01
MistralLoRA 15.15 13.54 13.48 12.10
VIKDFLlama 16.47 14.84 15.35 15.30
VIKDFQwen 16.27 14.65 15.22 15.29
VIKDFMistral 16.61 14.94 15.53 15.47
Furthermore, to assess the generalizability and robustness of VIKDF across different large language models, we integrated our framework with three state-of-the-art open-source LLMs: Llama-2, Qwen2.5 [51, 52], and Mistral-v0.3 [53], each with a parameter size of 7B. We first fine-tuned each LLM using LoRA [54] to establish baseline performances. Subsequently, we applied VIKDF to each LLM, following the same knowledge distillation and integration processes described earlier. The performance metrics, presented in Table 4, reveal that integrating VIKDF consistently enhances the BLEU-1 and ROUGE-L scores across all models and datasets. Notably, the improvements are more pronounced in the zero-resource conditions on the Image-Chat dataset. This significant enhancement underlines VIKDF’s effectiveness in leveraging implicit multimodal information to boost dialogue generation, especially when multimodal data is scarce. The consistent performance gains across different LLM architectures demonstrate the adaptability of our framework, showcasing its ability to seamlessly integrate with various models and improve their capabilities in challenging zero-resource scenarios. This underscores the robustness of VIKDF and its potential for broad applicability in advancing dialogue systems.
@@ 5.3 Ablation Study
Figure 6: Example of visual implicit knowledge textualized from dialogue context by VIKDF.
5.4 Case study
To provide a more detailed analysis of visual implicit knowledge, we attempt to textualize the visual implicit knowledge through the LLM, following the process illustrated in Figure 6. The process involves transforming the dialogue context $C$ through the IQ-Former, which distills visual implicit knowledge $K$ solely from $C$, and subsequently textualizing $K$ via the LLM using a text prompt. Importantly, no image input is provided at this stage; the model relies entirely on the textual dialogue context. We can see that the textual output illustrates the model’s capacity to encapsulate complex human emotions and curiosities about the sky, derived from the dialogue context. This demonstrates the model’s adeptness at integrating and expressing visual implicit knowledge. While the LLM possesses substantial world knowledge due to pre-training, the integration of the distilled knowledge $K$ allows it to capture visual implicit knowledge that is contextually relevant and may not be explicitly encoded in the LLM’s pre-trained weights. Thus, the LLM produces responses that reflect both its pre-existing knowledge and the newly distilled visual implicit knowledge pertinent to the dialogue context.
(a) Reddit Conversation
(b) Image-Chat (Zero Data)
@@ 5.4 Case study
5.5 Impact of Query Vector Quantity
@@
6 Conclusion and Future Work
@@
Acknowledgements
This research is supported by the National Natural Science Foundation of China (No. 62006034) and the Anhui Provincial Natural Science Foundation (2408085QF188).